OpenStudy (redd):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (anonymous):

assuming that ball travels vertically upwards we can use equations of constants acceo acceleration u = 42, g = acceleration due to gravity = 32, s = displacement = 0 s = ut - 0.5gt^2 0 = 42t - 16t^2 t(42-16t) = 0 t = 0 or 42/16 = 2.625 seconds approximately 2.6 seconds

OpenStudy (redd):

thank you so much

OpenStudy (anonymous):

ok